Show that 𝑦 = ln (π‘₯) is a solution of π‘₯𝑦′′ + 𝑦′ = 0Β 


1
Expert's answer
2021-11-05T09:36:35-0400

"xy" + y' = 0\\

so: y = ln(x) y' = \frac{1}{x y"} = \\frac{-1}{x^2}"

Plug this into the differential equation:



This shows that there is a solution, but not necessarily on what interval.

Now to show whic interval this is valid in, we need to use

the existence and uniqueness theorem.

Given a differential equation: "y" + p(x)y' = q(x) y(x0) = y_0"

A unique solution is guaranteed where p(x)0 and q(x) are continuous.



Put this into proper form by dividing through by x to get:



Thus:



These functions are both continuous on :